OpenStudy (anonymous):

At the end of the first day, 6 weeds appear in your neighborhood park. Each day, the number of weeds increases by four times. How many weeds will be in the park at the end of 14 days? Please show me how you get the answer. Thank you.

OpenStudy (anonymous):

lets say num of weeds at start = x at end of first day, weeds = x *4 end of 2nd day, x*4*4 = x*(4^2) 3rd x*4*4*4 = x*(4^3) end of 14th day = x*(4^14)

OpenStudy (cwrw238):

this is a geometric progression with a = 6 and common ratio = 4 you need sum of 14 terms = 6 * 4^14 - 1 ------- 4 - 1